the surface are of a cube is 726 square inches. what is the length of a lateral edge?
Answer by mananth(16946)

the surface are of a cube is 726 square inches. what is the length of a lateral edge?
surface area of cube = 6l^2
6l^2= 726
/6
l^2=121
l= 11 inches the side of the cube.
.
m.ananth@hotmail.ca